OpenStudy (anonymous):

medal (: if 9^x-9^x-1=24, then find the value of (6x)^x

OpenStudy (anonymous):

Okay so first solve this (9^x-9^(x-1) = 24) x = 3/2 ((1/2)*(6*3))^(3/2) = 27

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

Np :)

OpenStudy (anonymous):

xP